pearis:

Hihi please help!! Quadrilateral OPQR is inscribed inside a circle as shown below. What is the measure of angle O? You must show all work and calculations to receive credit. (10 points) Circle N is shown with an inscribed quadrilateral labeled OPQR. O is labeled 2x degrees. P is labeled y degrees. Q is labeled 2

Vocaloid:

Closed due to unanswerable (missing diagram)

ZombieLover:

opposite pairs of angles of an inscribed quadrilateral add to 180° O + Q = 180° 2x + 2x + 4 = 180° So now we can find Q. Q = 2x + 4, so plug in 44 for x to get Q = 2(44) + 4